In testing H0: µ = 3 versus Ha: µ ¹ 3 when =3.5, s = 2.5, and n = 100, what is the p-value?

a.0.0700

b.0.0228

c.0.0655

d.0.0456
Answer #1

For the given test scenario, we have

H0: µ = 3 versus Ha: µ ≠ 3

This is two tailed test.

Test statistic = Z = (Xbar - µ)/[S/sqrt(n)] = (3.5 – 3)/[2.5/sqrt(100)] = .5/.25 = 2

So, P-value = 0.0456

(by using z-table)

Answer: d.0.0456

[Note: P-value for one tailed test = 0.0456/2 = 0.0228]